### v3

# Deep Learning LOS Line-of-Sight

DL
Deep Learning
LOS
Line-of-Sight
NLOS
Non-Line-of-Sight
CH
Coverage Hole
UAV
Unmanned Aerial Vehicle
AGV
Automated Guided Vehicle
QoS
Quality of Service
CSI
Channel State Information
UE
User Equipment
CFRs
Channel frequency responses
RSS
Received Signal Strength
GBSM
Geometry-Based Stochastic channel Model
DL
Downlink
UL
Uplink
3GPP
3rd Generation Partnership Project
LoS
Line-of-Sight
ML
Machine Learning
LSTM
Long-Short Term Memory
TDD
Time-Division Duplex
OFDM
Orthogonal Frequency Division Multiplexing
LS
Least-Squares
SNR
Signal-To-Noise Ratio
MMSE
Minimum Mean Square Error
NMSE
Normalized Mean Square Error
MSE
Mean Square Error
QAM
Quadrature Amplitude Modulation
SSNR
Symbol-Signal-to-Noise Ratio
SNR
Signal-to-Noise Ratio
Adaptive 3D Geometry‐based Stochastic Channel Prediction for 3D DL Selection
Mervat Zarour1 Qiuheng Zhou1 Sergiy Melnyk1 and Hans D. Schotten12 2Department of Wireless Communication and Navigation, University of Kaiserslautern-Landau,
Kaiserslautern, Germany
schotten@rptu.de
1Intelligent Networks Research Group, German Research Center for Artificial Intelligence,
Kaiserslautern, Germany
{mervat.zarour, qiuheng.zhou, sergiy.melnyk, hans_dieter.schotten}@dfki.de
Abstract
This paper addresses the challenges of mobile user requirements in shadowing and multi-fading environments, focusing on the Downlink (DL) radio node selection based on Uplink (UL) channel estimation. One of the key issues tackled in this research is the prediction performance in scenarios where estimated channels are integrated. An adaptive deep learning approach is proposed to improve performance, offering a compelling alternative to traditional interpolation techniques for air-to-ground link selection on demand. Moreover, our study considers a 3D channel model, which provides a more realistic and accurate representation than 2D models, particularly in the context of 3D network node distributions. This consideration becomes crucial in addressing the complex multipath fading effects within geometric stochastic 3D 3GPP channel models in urban environments. Furthermore, our research emphasises the need for adaptive prediction mechanisms that carefully balance the trade-offs between DL link forecasted frequency response accuracy and the complexity requirements associated with estimation and prediction. This paper contributes to advancing 3D radio resource management by addressing these challenges, enabling more efficient and reliable communication for energy-constrained flying network nodes in dynamic environments.
Index Terms:
3D radio resource, link selection, channel prediction, LSTM, channel estimation, MMSE, LS, 3D channel model
publicationid: pubid: Copyright  © 2023 Your Name
I Introduction
Wireless communication networks are steadily gaining their popularity in all aspects of business and social life. Advances in wireless technologies, such as 5G and the upcoming 6G, promise to deliver faster data rates or reduced latency. Among others, these improvements will open up new use cases in the field of real-time applications, e. g. autonomous vehicles, wireless robotics control, and more. However, the critical challenge on the way to their realisation remains the reliability of the communication links. In outdoor scenarios, 2D terrestrial networks may be enhanced by the utilisation of the third coordinate by means of airborne wireless nodes.
3D radio resource nodes enable stable and highly reliable links to users, thanks to their high position flexibility within a 3D space. This capability addresses wireless signal propagation, blockage, reflection, shadowing, and the multipath effect. Compared to wireless links established through fixed terrestrial base stations, 3D radio resource nodes significantly improve wireless link quality.
The flight node can serve as base stations or relays and has high flexibility to improve the connectivity link quality with optimal 3D trajectory planning or altitude change, unlike the fixed terrestrial radio access node with a low Line-of-Sight (LoS) probability [1]. In 3D networking, ongoing research explores the advantages of employing over-the-air network nodes to enhance Quality of Service (QoS) while minimising transmission time due to energy limitation.[2]. In the mobility control use case, through consecutive control input via a wireless channel, the radio access node selection can support a highly stable and reliable link considering the DL ground-to-ground link prediction with high accuracy and energy constraint of the flight nodes [3]. In  [4], the author optimised the wireless service to the ground user using optimal cell partitioning, considering the average hover times of the Unmanned Aerial Vehicles.
Whereas utilising flight nodes is advantageous for maintaining a reliable communication link, the UAVs suffers from limited energy storage capability. To increase the energy efficiency of the flight nodes, we propose a 3D radio resource management system, which optimises the utilisation of power-demanding air-to-ground links based on QoS requirements. Thus, the ground-based users should preferably maintain the ground-to-ground link. In case of the ground link degradation due to i. e. fading, shadowing, or other effects, the air-to-ground link switch is performed preventing the violation of QoS requirements. Establishing an accurate 3D radio resource management system, relying on precise decision-making for DL adaptive transmission over-the-air links instead of ground-to-ground links is imperative [5]. So, the DL selection cannot be accurate when it is based on outdated channel information using uplink channel estimates due to the rapid variation in channel conditions, especially in dynamic environments such as urban areas with mobile User Equipment (UE).
Utilising Machine Learning (ML) methods for DL channel prediction enhances accuracy, mainly when the channel exhibits temporal correlation that influences the learning process of the predictor through interaction with channel state information. The channel predictor can capture the interaction between consecutive channel states in the time and frequency domain, thus benefiting from temporal channel correlation and enabling accurate channel prediction[6],[7]. Another important aspect is channel modelling, which incorporates 3D position instead of simplistic stochastic 2D channel modelling. A 3D channel model provides a more realistic perspective in the elevation angles (zenith) domain. This modelling approach included more parameter variables and their dependencies, which impact the channel variation [8]. [9] [10],[11].
To accomplish the task of optimal link selection, precise information on the condition of both ground-to-ground as well as air-to-ground links is required. In modern communication systems, the link condition is indicated by Mean Square Error (MSE). This measure is estimated based on the dedicated training signals sent in the UL channel. MSE is considered to be reciprocal, so the UL estimations is also utilised for DL communication. However, fast fading is known to destroy the reciprocity. Furthermore, especially real-time control applications may suffer the timeliness of the estimated MSE. The ability to predict DL channel behavior based on the UL estimation would significantly mitigate these effects.
This paper focuses on predicting Channel frequency responses (CFRs) in the context of 3D DL selection within an urban wireless signal propagation environment, utilising pilot-aided estimation methods and the 3D geometric stochastic 3GPP channel model considering the elevation domain. We analyse the prediction accuracy, considering the estimation error of the UL channel. Then, we introduce an innovative adaptive predictor structure for enhancing DL channel selection by leveraging upcoming channel information, thus ensuring mobile users’ highly reliable DL channel connection.
The paper is structured as follows. In the next section, the system model is presented. A detailed description of the simulation methodology follows in section III. Further, the results of our numerical analysis are discussed in detail in  IV. Finally, we summarise the main findings of this work in section V.
II System model
Figure. 1: System model
In an Urban Microcell environment, mobile network status changes dynamically, necessitating DL selection for optimising 3D radio resources and ensuring radio link quality through a closed-loop continuous monitoring and optimisation process. Air-to-ground DL usage occurs on demand to conserve flight node radio resources, guided by decisions from the ground-to-ground DL selector. The proposed framework, illustrated in figure 1, deals with radio channel propagation problems considering the 3D channel modelling. The wireless channel between transmitter and mobile UEs in 3D dimension e.g. Automated Guided Vehicles or UAVs, exhibits both slow fading due to path loss and shadowing effects and fast fading owing to multipath effects. This consideration considers the 3D positions of both the receiver and transmitter, providing a highly accurate description of signal propagation in a 3D network.
At the transmitter, the received pilot signal $P(k,t)$ in frequency block type serves to estimate the UL channel frequency response $\hat{H}_{UL}(k)$ for each subcarrier $k$ at the time $t$. The channel frequency responses ${H}_{DL}(k)$ for the upcoming DL are then interpolated linearly based on the estimated values. Our comparison explores the use of classical linear interpolation alongside deep learning-based DL channel prediction $\tilde{H}(k,t)$, built upon the estimation with two predictor structures.
In an open-loop fashion, the predictor updates the internal Long-Short Term Memory (LSTM) state based on the previously estimated value at each estimated subcarrier at time $t$. In the closed-loop scenario, the predictor forecasts the upcoming channel frequency response at each subcarrier based on the previous predicted value.
The proposed platform optimises the DL selector data using different channel estimation methods to improve the accuracy of the UL channel frequency responses. This is achieved using an adaptive predictor, which fits our computational demand and requirements and weighs the accuracy of the predicted DL data.
The data exchange between the transmitter and the receiver is performed in a framed manner. Each frame has a time interval of $dt$ for the UL and DL channel via the Time-Division Duplex (TDD) technique.
The estimated UL channel state information $\hat{H}$ can be outdated due to the fading effect when used at the selector input for the upcoming DL time slot. Then, the Link selector can serve the user with an air-ground DL when the ground-ground DL no longer meets the requirements.
The proposed channel model, estimator, and predictor are described in the following II-A,II-B,II-C.
II-A 3GPP 3D channel model
The 3GPP 3D channel modelling represents wireless channel propagation in a mobile user-scattering environment. We propose an urban microcell environment where the positions of the scatterers are not precisely determined but are characterised using stochastic channel parameters. We proposed a 3GPP 3D channel model to simulate the 3D ground-to-ground channel between a single transmitter and a mobile receiver[12].
Considering several key factors with its cross-correlation matrix, including shadowing effects, delay spread, azimuth, and zenith angle spread in both the arrival and departure directions. The number of considered channel taps is $L$. The channel response $h_{l}$ for the channel tap $l$ is represented as
$h_{l}=\sqrt{(P_{l})}.{(\gamma_{l}^{a},\theta_{l}^{a})}^{T}.\mathcal{F}.(\gamma_{l}^{d},\theta_{l}^{d}).\exp^{-j.2\pi.f.d}$ (1)
Where $P_{l}$ is the power of the $l^{th}$ path, and $d$ distance between $(x,y,z)$ transmitter and receiver positions, $f$ frequency, $(\gamma_{l}^{a},\theta_{l}^{a})$ arrival angles, and $(\gamma_{l}^{d},\theta_{l}^{d})$departure angles are determined via stochastic distributions based on measurements as detailed in  [13]. Where uncorrelated large scale parameter described as a vector of stochastic variables $X\sim\mathcal{N}(0,1)$: delay spread $X^{DS}$ and angel spread of departure ${X^{\theta}}^{d},{X^{\gamma}}^{d}$ and of arrive ${X^{\theta}}^{a},{X^{\gamma}}^{a}$. The shadowing effect via $X^{SF}$ and ricean $K$ factor via $X^{KF}$. The spatially exponential correlated vector X with de-correlation distance $d_{dec}$ means that the close-set UEs within $d_{dec}$ experience the same attention or path delay due to, e.g., common obstacle position. The path gain $P_{l}$ depends on the 3D distance, frequency, and the cross-correlated $X_{cor}=M.X$ where M is the cross-correlation matrix from  [13]. $\mathcal{F}$ is the coupling matrix to capture the polarisation change between the receiver and transmitter.
II-B Uplink channel estimator
Channel reciprocity in TDD Orthogonal Frequency Division Multiplexing (OFDM) system is precious in the adaptation of DL channel transmission. In the UL channel propagation duration, channel estimation is established to have an accurate channel response for each sub-carrier $k$ in OFDM system. Then, it can be based on channel reciprocity, assuming the channel response in the DL duration can interpolate it for the upcoming time slot [14]. Some decision-making for resource allocation, e.g. link selection, can be planning. The transmission adaption in DL gain depends on the channel variation behaviour and estimation accuracy. The channel frequency response is represented for each OFDM sub-carrier. The $H(k)$ is the true channel frequency response for sub-carrier $k$.
Least-Squares (LS) channel estimation estimates the frequency response of the wireless channel based on minimising the sum of squared errors between the received pilot symbols and the at-receiver known transmitted pilot symbols $P(k)$ for subcarrier $k$. It can be expressed as follows [15]:
$P_{r}(k)=P(k)\cdot H_{\text{UP}}+\sigma_{n}$ (2a)
$\hat{H}_{LS}(k)=P(k)^{-1}\cdot P_{r}(k)$ (2b)
LS Minimum Mean Square Error (MMSE) channel estimation finds the estimated channel response $\hat{H}(k)$ to minimise the mean-squared error of the estimation with considering the statistical property of the noise and the channel, so the LSMMSE estimation for the $k^{t}h$ subcarrier is $\hat{H}_{\text{LSMMSE}}$ at the pilot ${P(k)}$, can be as follows[16]:
$R_{hh}=\mathbb{E}\{{H_{\text{UP}}\cdot H_{\text{UP}}^{H}}\}$ (3a)
$\mathcal{W}=\frac{\beta\cdot{\sigma_{n}}^{2}}{\mathbb{E}\{P(k)^{2}\}}$ (3b)
$\hat{H}_{\text{LSMMSE}}(k)=R_{hh}\cdot(R_{hh}+\mathcal{W}\cdot{I})^{-1}\cdot\hat{H}(k)_{\text{LS}}$ (3c)
Where the $\mathbb{E}$ is the expectation value, $R_{hh}$ is the channel auto-correlation matrix, $\sigma_{n}$ the noise variation, and $\beta$ is constance dependent on signal constellation, in table  I.
II-C Downlink channel predictor
Based on the estimated UL channel response frequency $\hat{H}(k)$, can obtain an accurate, valuable channel response for adaptive transmission in the DL direction using a recurrent neural network, LSTM approach.
The accuracy of LSTM prediction depends on the purity of the input data to be used to update each memory cell $c^{i}(t)$ and the hidden state $h^{i}(t)$ for cell $i$ at time $t$ in the LSTM layer. Specifically, two approaches are to be established: First, the LSTM layer uses the current value from the estimator as input to update the memory cell and hidden state and then predicts the next value in the sequence.
Second, the LSTM predictor alternatively uses its own previous predicted value $\tilde{H}(k,t-1)$ as input to predict the channel response value for the upcoming time slot. So the LSTM cell has several components at each time $t$, including input gates $ig_{t}$, forget gates $fg_{t}$, output gates $og_{t}$, and the cell state $c_{t}$. These variables contribute to building and updating the hidden state for each LSTM cell $i$. So the LSTM recurrent layer output can be compressed as follows: Let the $x_{t}$ be the LSTM layer input from the sequence input layer.
$ig^{i}_{t}=\sigma(W_{xi}x_{t}+W_{hi}h_{t-1}+b_{i})$ (4a)
$fg^{i}_{t}=\sigma(W_{xf}x_{t}+W_{hf}h_{t-1}+b_{f})$ (4b)
$\tilde{c}^{i}_{t}=\tanh(W_{xg}x_{t}+W_{hg}h_{t-1}+b_{g})$ (4c)
$og^{i}_{t}=\sigma(W_{xo}x_{t}+W_{ho}h_{t-1}+b_{o})$ (4d)
$c^{i}_{t}=f_{t}\odot c_{t-1}+i_{t}\odot\tilde{c}^{i}_{t}$ (4e)
$h^{i}_{t}=o_{t}\odot\tanh(c^{i}_{t})$ (4f)
After the LSTM layer is a fully connected layer to benefit the LSTM layer output hidden state $h_{t}$. The full connected layer output is $h_{fc}$.
$h_{fc}=\sum_{i=1}^{N}w_{fc}^{j}\cdot h_{t}^{i}+b_{fc}$ (5)
Where $N$ is the number of LSTM cells, $W_{fc}^{i}$ is the weight $b_{fc}$ is the bias for the $c_{i}$ in the fully connected layer. The trained weight and bias $Wy,by$ in the regression layer to predict the upcoming channel response value aiming to minimise the loss function $L$ between the LSTM predicted and true value $x_{t}$.
$\tilde{H}(k,t+1)=Wy.h_{fc}+by$ (6a)
$L(\tilde{H}(k,t+1),x_{t})=\frac{1}{2}(\tilde{H}(k,t+1)-x_{t})^{2}$ (6b)
The normalised mean error Normalized Mean Square Error (NMSE) in 7 is used to represent the prediction accuracy.
$\text{NMSE}=\frac{1}{n\cdot m}\sum_{t=1}^{n}\sum_{k=1}^{m}(\frac{\tilde{H}_{dl}(k,t)-H_{dl}(k,t)}{\|{H_{dl}(k,t)}\|_{2}})^{2}$ (7)
The normalised predicted $\tilde{H}_{dl}(k,t)$ is compared to the actual $H_{dl}(k,t)$ to show the accuracy of the prediction, where $n$ is the number of predictions and $m$ is the number of OFDM subcarrier. Thus, the NMSE is calculated for both the real and imaginary parts of the predicted channel gain, ultimately obtaining the average NMSE.
III Simulated experimental scenario
We analysed an urban microcell scenario involving a single transmitter and a mobile user with a fixed speed along a linear path. We used OFDM modulation after Quadrature Amplitude Modulation (QAM) modulation. The Symbol-Signal-to-Noise Ratio (SSNR) is defined by $\frac{{E}\{P(k)^{2}\}}{\sigma^{2}_{n}}$ as shown in Equation 3b. And the Signal-to-Noise Ratio (SNR) is multiplied SSNR by the channel gain $|h|^{2}$. Decorrelation distance $d_{dec}$ is used to describe the similarity of channel variations, which depends on the distance between positions. The relative strength of the LoS component compared to the scattered components NLoS is stochastic variable $X^{kf}$.
Figure 2 illustrates how signals between a mobile device and the transmitter are affected in the urban environment. The obstacles can block, reflect, or scatter the wireless signals, creating multiple paths (channel taps) for the signal to reach the mobile user. The time axis is included to show how the channel gain varies rapidly due to the movement of the mobile user or other dynamic factors in the urban environment.
To build a robust deep-learning LSTM model, we divided our dataset into an $80\%$ training set and a test set. We explored two approaches: an open-loop method based on previously estimated values and a closed-loop approach based on previously predicted values. The training and testing procedures were repeated several times with different generated channel state information, using the simulation setup described in table I.
We evaluated the performance of the model by calculating the average of its predictions. The performance metric we used is the average of all predicted values for multiple subcarriers, as shown in the equation7.
This comprehensive analysis allowed us to assess the effectiveness and robustness of the model under multi-generated CSI under the same scenario conditions.
Figure. 2: Channel gain with 20 channel taps, mobile user in urban microcell scenario
TABLE I: Simulation Setup Parameters
Parameter Value
OFDM DFT size 128
Number of Channel taps $L$ 5
UE path-length 100 m
UE speed 10 m/s
Channel sampling time 5 ms
Decorrelation distance $d_{dec}$ 5 m
$X^{KF}$ parameter $\mu_{KF}=-3$\mathrm{d}\mathrm{B}$$, $\sigma_{KF}=0.5$\mathrm{d}\mathrm{B}$$
CSI size 2000
Data set size 20000
Modulation 16QAM
$\beta$ 17/9 [16]
Training algorithm adam
Number of Hidden neurons $N$ 200
IV Result
After the estimation, we have two options to provide the DL selector. Either we still interpolate the estimated channel frequency responses linearly or provide it to the channel predictor. The received pilot symbols in a frequency block manner offer the possibility of using an estimation method to estimate the channel’s frequency response at each OFDM sub-carrier. The LS estimator provides a low-complexity method compared to the MMSE method due to the matrix inversion and the requirement to acknowledge the auto-correlation matrix of the channel responses.
As illustrated in Figure 3, the MMSE estimator performs better in bad channel conditions when the SNR is low, as it minimises the mean square error after linear interpolation in the DL based on the pilot symbols via the UL channel, in comparison to the LS method. Both methods converge to the ideal estimation following interpolation under good channel conditions.
Based on these results, we analyse the adaptive predictor, which has two modes for predicting the upcoming channel frequency responses, as shown in Figure 4. Using the best estimation method, we compare the previous results in Figure 3 with our two proposed predictor modes. Despite the estimation error, the open-loop predictor provides highly accurate DL channel frequency responses. This means that we should provide the predictor with highly accurate estimated UL channel information each time slot.
Compared to the closed-loop predictor, which has lower accuracy than the open-loop, we do not need the estimated value for prediction but instead rely on the previously predicted values for the next output. This results in lower complexity compared to the open-loop predictor mode. Therefore, when using MMSE estimation, there is a trade-off between complexity, computation, estimation frequency, and the accuracy of the input data for the link selector.
Figure. 3: DL channel information correctness in case of using linear interpolation with various estimation methods
Figure. 4: DL channel information correctness in case of use prediction with high accurate estimation method
In Figure 5, we analyse the adaptive predictor in the case of a user with low-accuracy LS estimation. Closed-loop prediction exhibits superior accuracy compared to the interpolation method. Therefore, closed-loop prediction mitigates complexity, particularly at low estimation, and enhances overall accuracy when applying LS estimation.
Figure. 5: DL channel information correctness in case of use prediction with low accurate estimation method
A significant challenge arises in the context of open-loop prediction utilising LS estimation due to the accumulation of estimation errors at the predicted output. This accumulation is a consequence of consecutive state updates within LSTM layers. Consequently, it leads to diminished accuracy in estimations and a marked reduction in prediction performance when contrasted with our baseline approach involving linear interpolation.
Hence, it is evident that both prediction modes hold the potential to generate output based on DL selector input while considering long-term dependencies. However, the adaptability of the prediction system we propose is paramount. It enables the benefit of predictions based on the chosen estimation method and the specific accuracy requirements for the input data. This adaptability ensures the versatility of our system, enabling it to meet varied system complexity and deliver the desired level of precision.
So, utilising closed-loop prediction results in heightened accuracy and reduced complexity, particularly at lower estimation frequencies, while open-loop prediction coupled with LS estimation may confront challenges associated with error accumulation.
V Conclusion
This paper presents a deep learning LSTM approach for predicting frequency-domain channel behaviour for a 3D stochastic fading channel within an urban environment, 3GPP geometric channel model. Our approach considers various factors, including user mobility, shadowing, and multipath effects.
The proposed method involves adaptive channel prediction based on pilot-aided estimation. We assess the accuracy of the predicted channel frequency responses, which serve as input to the 3D DL selector decision-making process.
We proceed to analyse the prediction gain between two prediction possibilities. The first option involves making predictions after each estimation in an open-loop fashion, while the second entails employing consecutive closed-loop predictions for improved accuracy. This choice represents a trade-off between estimation requirements and prediction accuracy. In case of estimation accuracy is high, consecutive closed-loop predictions can yield superior results. However, the closed-form prediction strategy may lose accuracy if the LS estimation accumulates estimation errors due to long-term dependencies.
The adaptive prediction strategy can balance estimation and prediction complexity and the required accuracy at the input of the 3D DL selector.
Acknowledgements
The authors acknowledge the financial support by the Federal Ministry of Education and Research of the Federal Republic of Germany (BMBF) in the project “Open6GHub” with funding number (DFKI-16KISK003K). The authors alone are responsible for the content of the paper.